Grade Select GradePhysical Chemistry

loweringnof vapou pressure due to 1 molal aqueous sotion at 1000c is....

Asokkumar Verma
10 Years agoGrade Select Grade

1 Answer

Beenu Mathew
10 Years ago
Hi,
 
P0a – Pa/P0a = mole fraction of solute
(we all know this fact)
Now 
you have been given molality= 1 mole/kg
Now this thing above can be changed to 1 mole of solute in 1000g/1kg solvent(which is the definition of molality)
mole fraction of solute =  1/1+ 1000/18 (solvent is water )
= 1/56.5
NOw, phew!! we got this done
now what was it you wanted lowering of vapor pressure
ieP0a – Pa = P0a* mole fraction of solute
                 = 760 torr * 1/56.5
                 =13.4 torr
 
THANKS 
BEST OF LUCK!!
PLEASE APPROVE!!
:D